OpenStudy (briana.img):

One student begins in Cleveland, Ohio, and walks towards Buffalo, New York, a distance of 200 miles. She walks at the rate of 2.5 miles per hour. Another student begins riding his bicycle in Cleveland and rides towards Buffalo at the rate of 16 miles per hour. The walking student begins her journey at 8:00 AM. The riding student starts 4 hours later.

OpenStudy (briana.img):

Part A Write a system of two equations that could be used to find the distance each student is from Cleveland, Ohio, at any given time. Let x represent the distance that the walking student is from Cleveland, let y represent the distance that the riding student is from Cleveland, and let t be the time that has elapsed since the walking student left Cleveland. Part B Solve the system of equations for x and y to determine how many hours elapse before the students meet.

OpenStudy (briana.img):

What method is used for this? What is the equation?

OpenStudy (anonymous):

lot of rules for this one i am sure we can do it

OpenStudy (briana.img):

@satellite73 yeah have no idea what method to use or how the equation would even look like

OpenStudy (anonymous):

i think mostly it is distance is rate times time

OpenStudy (anonymous):

She walks at the rate of 2.5 miles per hour her rate is \(2.5\) so if her time is \(t\) then her distance is \(2.5t\)

OpenStudy (anonymous):

similarly, the riding students distance is \(16t\) BUT the walker has a four hour head start

OpenStudy (anonymous):

she walked for four hours before the riding student started in four hours she went \(2.5\times 4=10\) miles

OpenStudy (anonymous):

so add on 10 miles to her distance to make it \[2.5t+10\] and the rider is \[16t\] and evidently they meet when their distances are the same set \[2.5t+10=16t\] and solve for \(t\)

OpenStudy (briana.img):

@satellite73 but doesn't there need to be an x and y?!

OpenStudy (briana.img):

and two equations?

OpenStudy (anonymous):

ok really i object because it makes it sound like god give us one way to solve a problem i really don't know how to do it with an \(x\) and \(y\) but the above answer i assure you is correct we can try to bend to their method, but i am not sure how to do it in a natural way

OpenStudy (briana.img):

@satellite73 sorry i was just going with what the problem was asking

OpenStudy (anonymous):

x represent the distance that the walking student is from Cleveland ok lets put \[x=2.5t+10\]

OpenStudy (anonymous):

and put \[y=16t\]

OpenStudy (briana.img):

@satellite73 yeah that makes more sense now

OpenStudy (anonymous):

and since they have to be equal (as they meet at the same distance) set \[2.5t+10=16t\] and solve for \(t\)

OpenStudy (anonymous):

you good from there?

OpenStudy (briana.img):

@satellite73 i think so..? just trying to figure out how i would solve it

OpenStudy (anonymous):

subtract \(2.5t\) from both sides

OpenStudy (anonymous):

then divide

OpenStudy (briana.img):

10/13.5t ?!!?

OpenStudy (anonymous):

i.e. \[t=\frac{10}{13.5}\]

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (briana.img):

so t = 1.35 ?!

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{13.5}{10}=1.35\] this is upside down

OpenStudy (anonymous):

\[\frac{10}{13.5}=\frac{100}{135}\] which i guess you can reduce or turn in to a decimal

OpenStudy (briana.img):

2.5(0.740)+10 = 11.85 16(0.740) = 11.84

OpenStudy (anonymous):

yeah i got a time of \(.74\) hours

OpenStudy (briana.img):

@satellite73 thank you!!

OpenStudy (anonymous):

yw